# Unconstrained Open Vocabulary Image Classification: Zero-Shot Transfer from Text to Image via CLIP Inversion

### v2

@@ 5 Conclusion
In this paper, we address the problem of unconstrained real-time open vocabulary image classification, and highlight the limitations of current state-of-the-art methods in accomplishing this task. To overcome these limitations, we propose an autoregressive object decoder trained on a large-scale text-only dataset. The decoder generates object nouns directly as language on the basis of CLIP embeddings, and crucially, does not require any class candidates to be provided. To create the text-only training dataset, we introduce a comprehensive object noun dictionary and employ multiset prompt-templating in combination with LLM-generated captions and a pivotal noise augmentation strategy to effectively bridge the modality gap from text to images. Our approach demonstrates promising results on both established and newly proposed in-the-wild classification datasets. We lift the common limitation of zero-shot image recognition models regarding prompts and class candidates, ultimately allowing agents to immediately recognize and label any object in dynamic environments. Future work will include providing an input to the decoder that controls the level of semantic granularity used in the predicted output nouns.
Ethical AI.
@@ A.3 Variance-Balanced Transformer Initialization
  • Input stage: A linear projection of the input embedding vector in parallel to the token embeddings of the ground truth object noun tokens, followed by an additive positional encoding,
@@
To ensure approximate unit scale throughout the transformer post-initialization, we wish for each signal element throughout the network to have near-unit variance. As such, we initialize the weights $W_{p}$ of the linear projection layer from a zero-mean normal distribution of variance $\tfrac{1}{2}$, i.e. $\mathcal{N}(0,\tfrac{1}{2})$, and do the same for all weights in the token embeddings $W_{te}$ and learned position encodings $W_{pe}$. If $W_{te}^{*}$ is the learnable token embeddings for the specific ground truth tokens being passed to the object decoder, then from Eqs. 1, 2 and 3, we can see that the expected variance $\sigma_{p}^{2}$ post-positional encoding, i.e. of the actual input sequence vectors to the transformer layers, is thus
@@ Appendix C LLM-based Caption Generation
Given the complete object noun dictionary, an LLM was used to generate rich dynamic captions involving each of the object nouns in the dictionary. Based on the log-frequencies obtained from the object noun dictionary, ten times as many captions are generated independently for singular and plural forms of the noun, with at most 100 being generated for either category. OpenAI GPT-3.5 Turbo was used, specifically gpt-3.5-turbo-1106 with a temperature of 0.8, maximum 1000 output tokens, and example prompts/responses as shown in Appendix C. Ten captions are generated in every API call, and parsed using all available spelling variations of the object noun into appropriate caption templates. For instance, ”Two children playing with a bicycle in the yard” is parsed and stored as ”Two children playing with a {singular} in the yard”. In cases of words where there are multiple spellings, like for example orangutan, this ensures that a caption template can be used in future to generate data for any desired spelling if required by the sampling strategy. A total of 1.8M captions were generated in this way.
@@ Appendix D Effect of Noise Augmentation
The standard deviation of the angle $\beta$ can empirically be measured to be 1.94. This narrow distribution of $\beta$, i.e. the distribution of angle separations that occur when using the nominal Gaussian noise augmentation strategy, is shown in Fig. 5. Also shown in the plot is the measured distribution of angle separations for matching and non-matching image-label pairs of the ImageNet-1K validation set, i.e. demonstrating the modality gap as well as the very thin separation that CLIP models provide between matching and non-matching pairs. The positive distribution-widening effect of the nominal strategy of adding 15% uniform angle noise from 45–75 is also shown.
@@ Appendix F Multimodal LLM Image Annotation
Due to the recent availability of some multimodal LLMs with a very high level of image understanding, it is conceivable to approximately quantify the performance of open vocabulary classification models by auto-annotating their predictions on image sets using these LLMs. This is particular useful when it is just not feasible to perform the required amount of annotations as a human. We use OpenAI GPT-4o, specifically gpt-4o-2024-05-13 with a temperature of 0.2, maximum 512 output tokens, and a cumulative probability threshold of 0.6 for nucleus sampling. The chosen parameters reflect the stability and consistency that is desired from the LLM. The prompt used for auto-annotation is shown in Appendix F for an example image. Asking the LLM to first provide a description of the image allows it to adjust its summarized knowledge of the image according to the provided object noun candidates, and enhances its classification precision. For best results, 8–15 object noun candidates are always given, as providing too few or too many has a small potential of causing inconsistent annotations for fine-grained object nouns. With intentional randomization of candidate ordering and which candidate is picked for which API request, each object noun is annotated 3–5 times, stopping before 5 if there is at least 80% agreement in the LLM’s opinions. If the final agreement is at least 80%, then the corresponding opinion (correct or incorrect) is taken as the final annotation, otherwise the annotation is close. No primary vs. secondary classification is made because it was experimentally deemed too inconsistent. Due to the relatively unlikely condition required for the close category (LLMs tend to be wrong more often than they are inconsistent at a low temperature), it is systematically not seen as often as for human annotations.
@@ Appendix G Model Hyperparameters
Configuration World-H Wiki-L IN1K
Prefix tokens $P=2$ 77.91 62.75 65.51
Prefix tokens $P=4$* 78.92 63.97 68.11
Prefix tokens $P=6$ 78.03 64.15 68.08
Layers $L=4$ 77.23 62.38 63.85
Layers $L=6$* 78.92 63.97 68.11
Layers $L=8$ 77.94 62.78 69.05
Layers $L=12$ 78.15 64.00 71.19
Hidden dim $H=384$ 76.97 61.65 65.77
Hidden dim $H=512$* 78.92 63.97 68.11
Batch size $B=4096$ 78.14 63.73 66.76
Batch size $B=8192$* 78.92 63.97 68.11
Batch size $B=16384$ 79.56 63.17 67.85
Epochs $E=12$ 77.33 63.30 66.33
Epochs $E=18$* 78.92 63.97 68.11
Epochs $E=24$ 77.77 63.00 67.71
Table 6: Object decoder architecture and training hyperparameter ablations (mean of 3). IN1K is the ImageNet-1K classification performance of the trained decoders, and the asterisks correspond to the nominal ablation baseline (FT2). World-H and Wiki-L are prediction scores (described in Sec. 4) on the corresponding datasets. See Sec. 4.2 for more ablation results.
@@ Appendix I Decoder Architecture Ablations
Ablations of the core NOVIC model hyperparameters are shown in Tab. 6. The results support the nominal choice of hyperparameters, and suggest that slight gains could potentially be made with a greater number of transformer layers (although model generalization becomes slightly less reliable) and/or batch size. Refer to Sec. 4.2 for ablation results pertaining to the training dataset, noise augmentation, and CLIP model.

### v3

@@ 5 Conclusion
In this paper, we address the problem of unconstrained real-time open vocabulary image classification, and highlight the limitations of current state-of-the-art methods in accomplishing this task. To overcome these limitations, we propose an autoregressive object decoder trained on a large-scale text-only dataset. The decoder generates object nouns directly as language on the basis of CLIP embeddings, and crucially, does not require any class candidates to be provided. To create the text-only training dataset, we introduce a comprehensive object noun dictionary and employ multiset prompt-templating in combination with LLM-generated captions and a pivotal noise augmentation strategy to effectively bridge the modality gap from text to images. Our approach demonstrates promising results on both established and newly proposed in-the-wild classification datasets. We lift the common limitation of zero-shot image recognition models regarding prompts and class candidates, ultimately allowing agents to immediately recognize and label any object in dynamic environments. Future work can include providing an input to the decoder that controls the level of semantic granularity used in the predicted output nouns.
@@ A.3 Variance-Balanced Transformer Initialization
  • Input stage: A linear projection of the input embedding vector in parallel to the token embeddings of the ground truth object noun tokens, followed by an additive positional embedding,
@@
To ensure approximate unit scale throughout the transformer post-initialization, we wish for each signal element throughout the network to have near-unit variance. As such, we initialize the weights $W_{p}$ of the linear projection layer from a zero-mean normal distribution of variance $\tfrac{1}{2}$, i.e. $\mathcal{N}(0,\tfrac{1}{2})$, and do the same for all weights in the token embeddings $W_{te}$ and learned position embeddings $W_{pe}$. If $W_{te}^{*}$ is the learnable token embeddings for the specific ground truth tokens being passed to the object decoder, then from Eqs. 1, 2 and 3, we can see that the expected variance $\sigma_{p}^{2}$ post-positional embedding, i.e. of the actual input sequence vectors to the transformer layers, is thus
@@ Appendix C LLM-based Caption Generation
Given the complete object noun dictionary, an LLM was used to generate rich dynamic captions involving each of the object nouns in the dictionary. Based on the log-frequencies obtained from the object noun dictionary, ten times as many captions are generated independently for singular and plural forms of the noun, with at most 100 being generated for either category. OpenAI GPT-3.5 Turbo was used, specifically gpt-3.5-turbo-1106 with a temperature of 0.8, maximum 1000 output tokens, and example prompts/responses as shown in Appendix C. Ten captions are generated in every API call, and parsed using all available spelling variations of the object noun into appropriate caption templates. For instance, "Two children playing with a bicycle in the yard" is parsed and stored as "Two children playing with a {singular} in the yard". In cases of words where there are multiple spellings, like for example orangutan, this ensures that a caption template can be used in future to generate data for any desired spelling if required by the sampling strategy. A total of 1.8M captions were generated in this way.
@@ Appendix D Effect of Noise Augmentation
The standard deviation of the angle $\beta$ can empirically be measured to be 1.94. This narrow distribution of $\beta$, i.e. the distribution of angle separations that occur when using the nominal Gaussian noise augmentation strategy, is shown in Fig. 5. Also shown in the plot is the measured distribution of angle separations for matching and non-matching image-label pairs of the ImageNet-1K validation set, i.e. demonstrating the modality gap as well as the very thin separation that CLIP models provide between matching and non-matching pairs. The positive distribution-widening effect of the nominal strategy of adding 15% uniform angle noise from 45–75 is also shown (labeled “Gauss 3.25 $+$ Uniform” in Fig. 5). With this nominal strategy, for 85% of all samples Gaussian noise of magnitude $G=3.25$ is added, and for the remaining 15% of samples (randomly chosen), uniform angle noise of random magnitude 45–75 is applied.
@@ Appendix F Multimodal LLM Image Annotation
Due to the recent availability of some multimodal LLMs with a very high level of image understanding, it is conceivable to approximately quantify the performance of open vocabulary classification models by auto-annotating their predictions on image sets using these LLMs. This is particularly useful when it is just not feasible to perform the required amount of annotations as a human. The general process is as follows:
  1. 1.
    The chosen to-be-evaluated model(s) are inferenced on the required dataset of images, and the predicted object nouns are recorded (saving the top-k predictions for each image is also possible).
  2. 2.
    For each individual image, the deduplicated set of all predicted object nouns across all models is collected.
  3. 3.
    Groups of up to 15 of these predicted object nouns are evaluated per call to the LLM, until every object noun has up to 5 opinions from the LLM whether it is correct or incorrect (present in the image or not).
  4. 4.
    A ground truth annotation is established per-image per-predicted object noun based on the opinions that were given by the LLM.
@@
We use OpenAI GPT-4o, specifically gpt-4o-2024-05-13 with a temperature of 0.2, maximum 512 output tokens, and a cumulative probability threshold of 0.6 for nucleus sampling. The chosen parameters reflect the stability and consistency that is desired from the LLM. The prompt used for auto-annotation is shown in Appendix F for an example image. Asking the LLM to first provide a description of the image allows it to adjust its summarized knowledge of the image according to the provided object noun candidates, and enhances its classification precision. For best results, 8–15 object noun candidates are always given, as providing too few or too many has a small potential of causing inconsistent annotations for fine-grained object nouns. With intentional randomization of candidate ordering and which candidate is picked for which API request, each object noun is annotated 3–5 times, stopping before 5 if there is at least 80% agreement in the LLM’s opinions. If the final agreement is at least 80%, then the corresponding opinion (correct or incorrect) is taken as the final annotation, otherwise the annotation is close. No primary vs. secondary classification is made because it was experimentally deemed too inconsistent. Due to the relatively unlikely condition required for the close category (LLMs tend to be wrong more often than they are inconsistent at a low temperature), it is systematically not seen as often as for human annotations.
@@ Appendix I Decoder Architecture Ablations
Ablations of the core NOVIC model hyperparameters are shown in LABEL:tab:decoder_ablations. The results support the nominal choice of hyperparameters, and suggest that slight gains could potentially be made with a greater number of transformer layers (although model generalization becomes slightly less reliable) and/or batch size. Refer to Sec. 4.2 for ablation results pertaining to the training dataset, noise augmentation, and CLIP model.
Appendix J Image Classification Benchmark Trends
Tab. 1 shows the performance of NOVIC on various image classification benchmarks, many of which are commonly used to benchmark the zero-shot image classification performance of CLIP models. In general, the trend can be seen that the classification performance of NOVIC models slowly drops off as the complexity and extensiveness of the training data is increased towards the full size of the object noun dictionary. This can be expected, as the object decoder has increasingly many object concepts to learn as the frequency threshold (FT) decreases, and these concepts are also increasingly densely packed in the embedding space, making them harder to accurately keep apart. An exception to this general trend for some of the classification benchmarks is the FT9 score, as the FT9 training data does not actually always cover all of the target nouns required for each dataset. This concretely means that $412/1000$ ImageNet-1K classes, $17/200$ Tiny ImageNet classes, $50/200$ ImageNet-A classes, and $37/200$ ImageNet-R classes, cannot contribute to the final classification score, resulting in the disadvantaged lower scores marked with a dagger ($\dagger$) in Tab. 1. Another exception to this general trend are the benchmarks that have few and coarse-grained classes, such as Imagenette and CIFAR-10. For these datasets, the performance depends less on the diversity of object nouns seen during training, and more on just the raw ability of the CLIP image encoder to produce an accurate and representative embedding, resulting in fairly consistent classification performances irrespective of the frequency threshold.
Model Prompt-free Classification Text-only Training Seen Objects Used Objects RAM Vocab Primary Score Prediction Score Overall Score
Tag2Text [29] 3 429 176 91.9 62.32 ${\scriptstyle\xleftarrow{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-15.29}}}$ 77.61 ${\scriptstyle\xrightarrow{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-13.49}}}$ 64.12
RAM [76] 4 585 160 100.0 57.35 ${\scriptstyle\xleftarrow{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-17.65}}}$ 75.00 ${\scriptstyle\xrightarrow{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-15.86}}}$ 59.14
RAM (FT2) 11 897 188 76.8 64.71 ${\scriptstyle\xleftarrow{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-11.61}}}$ 76.32 ${\scriptstyle\xrightarrow{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-13.84}}}$ 62.48
RAM (FT0) 42 919 188 74.3 63.97 ${\scriptstyle\xleftarrow{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-11.03}}}$ 75.00 ${\scriptstyle\xrightarrow{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-13.84}}}$ 61.16
NOVIC (FT9) 2 919 235 66.9 74.82 ${\scriptstyle\xleftarrow{{\color[rgb]{1,0.5,0.3}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.5,0.3}\pgfsys@color@cmyk@stroke{0}{0.50}{0.70}{0}\pgfsys@color@cmyk@fill{0}{0.50}{0.70}{0}-2.79}}}$ 77.61 ${\scriptstyle\xrightarrow{{\color[rgb]{1,0.5,0.3}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.5,0.3}\pgfsys@color@cmyk@stroke{0}{0.50}{0.70}{0}\pgfsys@color@cmyk@fill{0}{0.50}{0.70}{0}-3.22}}}$ 74.39
NOVIC (FT6) 5 899 252 58.5 76.47 ${\scriptstyle\xleftarrow{{\color[rgb]{1,0.5,0.3}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.5,0.3}\pgfsys@color@cmyk@stroke{0}{0.50}{0.70}{0}\pgfsys@color@cmyk@fill{0}{0.50}{0.70}{0}-2.21}}}$ 78.68 ${\scriptstyle\xrightarrow{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}-1.01}}}$ 77.67
NOVIC (FT2) 11 897 253 52.6 77.94 ${\scriptstyle\xleftarrow{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}-1.62}}}$ 79.56 ${\scriptstyle\xrightarrow{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}-0.64}}}$ 78.92
NOVIC (FT0) 42 919 251 49.6 71.14 ${\scriptstyle\xleftarrow{{\color[rgb]{1,0.5,0.3}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.5,0.3}\pgfsys@color@cmyk@stroke{0}{0.50}{0.70}{0}\pgfsys@color@cmyk@fill{0}{0.50}{0.70}{0}-3.53}}}$ 74.67 ${\scriptstyle\xrightarrow{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}-1.10}}}$ 73.57
NOVIC (FT2) 11 897 263 55.5 86.77 ${\scriptstyle\xleftarrow{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}-1.17}}}$ 87.94 ${\scriptstyle\xrightarrow{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}-0.81}}}$ 87.13
NOVIC (FT0) 42 919 260 48.5 86.95 ${\scriptstyle\xleftarrow{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}-1.32}}}$ 88.27 ${\scriptstyle\xrightarrow{{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}-0.37}}}$ 87.90
Table 6: Open vocabulary World-H results in comparison to baselines (best of 3). Tag2Text and RAM achieve artificially high scores by frequently outputting uninformative generic tags like man, tree, grass, plate, and often miss the actual intent of an image. Scores are in %. RAM Vocab is the percent of samples whose predictions are in the core RAM vocabulary. Primary Score is the Prediction Score when rewarding only primary predictions. Overall Score is the Prediction Score with $\times$0.5 applied to the score contributions of any coarse predictions, effectively penalizing vague generic predictions. See Tab. 3 for Wiki-H results. Using DFN-5B H/14-378 CLIP model.
Training on large-scale datasets is also inherently not deterministic—in particular also due to the randomized data (e.g. noise) augmentation—and some natural variations in the results will occur even if taking the mean of three training runs. The object decoder has thousands to tens of thousands of concepts to learn, and evaluating on a classification dataset with e.g. 10–101 classes is probing a small fraction of what the model has learned and is rewarded for by its training loss. This means that two models with identical loss will always have some concepts that they have learned better, and some they have not learned quite as robustly. If that happens to align with the classes required for a classification dataset, the models then either do slightly better or slightly worse on them. As previously mentioned, CIFAR-10 is an example of all models doing about the same, up to natural variation, as the dataset classes are very few and coarse, and the dataset is relatively ‘easy’.
The classification results for the FT6 and FT2 NOVIC models are similar for many of the benchmarks, as both training sets have a reasonable size and comfortably cover all required class names as target nouns. The dip in performance for FT0 is more pronounced in general, due to the near-quadrupling relative to FT2 of the number of object concepts to learn, with the magnitude of the dip being tendentially correlated to larger numbers of classes and more fine-grained classification concepts overall. Both of these factors result in smaller regions of the embedding space, with less cosine distance between them, needing to be kept apart. This, in combination with the more finely ‘concept-subdivided’ embedding space of an FT0 model, leads to the performance dip.
Appendix K Additional Baseline Comparisons
In Tabs. 3 and 6, we compare (as described in Sec. 4.1) NOVIC to the image tagging baselines Tag2Text [29] and RAM [76], on the Wiki-H and World-H datasets respectively. While the image tagging baselines require a list of noun candidates to be supplied at inference time and are trained on image-text pairs, NOVIC performs completely prompt-free image classification and is trained much more efficiently on text-only data. Also, while NOVIC is trained primarily on a synthetic text dataset constructed in a controllable and unbiased manner from an entire English dictionary of object nouns, the baselines train on ‘only’ 14M image-text pairs, which have a significantly lower diversity of seen objects, and do not offer any controllability of the distributions of the nouns and image embeddings seen during training. Furthermore, Tag2Text has a fixed set of labels, while RAM attempts to provide some level of open vocabulary recognition beyond its nominal set of labels by distilling an image encoder from the base CLIP image encoder. Due to the limited size of the used image-text dataset however, and noting that the core set of labels was already constructed by parsing these exact text captions, one can understand that the ‘open’ abilities of RAM are somewhat limited, i.e. RAM is not truly open vocabulary in any way close to how one considers CLIP to exhibit open vocabulary abilities.
In Tabs. 3 and 6, we can see that NOVIC has seen significantly more diverse object nouns during training (42 919 as opposed to less than a few thousand for the baselines), and this is clearly reflected by its ability to use a wide array of object nouns in its predictions—generally close to as many unique nouns as there are images in the corresponding dataset (e.g. 263 unique predictions for 272 images). The tendency of the baselines not to provide such a diverse set of predictions can be observed from their significantly lower number of used objects. The restricted ability of RAM to perform open vocabulary classifications can also be identified by observing the high proportion of its generated predictions that belong to the core RAM vocabulary, even when provided with a large variety of object nouns to choose from (e.g. FT0). RAM’s open vocabulary predictions are also observed to have a significantly higher error rate than core vocabulary predictions.
Three prediction scores are calculated for each model—Prediction Score, the standard prediction score as defined in Sec. 4, Primary Score, which does not reward secondary predictions and thereby emphasizes the important ability of the model to focus on the actual intended object of an image, [^9] Note that even the primary score does not adequately penalize the baselines’ scores in practice however, for example in instances where a plate of food is the focus of the image and the baselines invariably predict plate whereas NOVIC tends to predict the specific name of the dish. and Overall Score, which adjusts the standard prediction score to reduce the reward for overly coarse predictions and thereby emphasizes the important ability of the model to provide meaningful and informative fine-grained classifications. It should be noted that both the primary and overall scores can by definition only at most be equal to the standard prediction score, so an ideal model that emits only fine-grained primary predictions would have all three scores be equal. We can observe that for NOVIC this is indeed almost the case, whereas the baselines exhibit very significant drops in score when penalizing secondary and coarse predictions as such. Overall, this demonstrates that NOVIC produces high quality predictions that are simultaneously diverse and fine-grained yet at the same time accurate, as well as correctly focusing on the main object of an image, all while being real-time, prompt-free, truly open vocabulary, and very efficiently trained with text only. This makes NOVIC a strong improvement over the state-of-the-art baselines.
Appendix L Ethical AI